OpenStudy (anonymous):

Which of the following graphs is the same as y = log1/2x? a. y = 2x b. y = 2-x c. y = -(2x) I really don't understand logarithms with fractions and decimals

OpenStudy (anonymous):

49x2 – 70x + 25 A. The coefficient does not permit the factoring. B. The coefficient does permit the factoring, but the x coefficient does not permit the factoring. C. The constant value does not permit the factoring. D. The polynomial may be factored into a perfect square.

OpenStudy (anonymous):

log1/2x= log1-log2x log1=0 so y=log-2x

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Which of the following is a factor of the polynomial below? A. x – 2 B. x + 2 C. x + 5 D. x + 8

OpenStudy (anonymous):

You didn't even post a polynomail

OpenStudy (anonymous):

where is polinomial??

OpenStudy (anonymous):

love answer of your first question is C

OpenStudy (anonymous):

it didnt show one second

OpenStudy (anonymous):

Monica is attempting to factor the expression 18x2 + 27x – 5. In her first step, she rewrites the expression as 18x2 + 32x – 5x – 5. Which of the following statements is true about Monica's work? A. The first step is correct. She should continue to factor the expression. B. The two middle coefficients add to the correct number, but do not multiply to the correct number. C. The two middle coefficients multiply to the correct number, but do not add to the correct number. D. The two middle coefficients do not add or multiply to the correct number

OpenStudy (anonymous):

B correct middle expression answer would be 30x-3x

OpenStudy (anonymous):

Which of the following strategies is appropriate for factoring polynomials with 2 terms? A. determine if the polynomial is the sum of two squares B. determine if the polynomial is the difference of two squares C. determine if the polynomial is a perfect square trinomial D. determine if it can be factored by grouping